OpenStudy (anonymous):

You invest $3,000 in an account with an interest rate of 5.5% compounded continuously. How much money would be in the account after 5 years?

OpenStudy (slaaibak):

3000 * e^nr

jimthompson5910 (jim_thompson5910):

In this case, P = 3000, r =0.055 and t = 5 We're going back to the formula A = Pe^(rt) because we're compounding continously

jimthompson5910 (jim_thompson5910):

Plug these values in to get A = 3000*e^(0.055*5)

OpenStudy (anonymous):

A= 300*e^0.275

jimthompson5910 (jim_thompson5910):

good so far

OpenStudy (anonymous):

Sorry A=3000*e^0.275

jimthompson5910 (jim_thompson5910):

oh good point, I glossed over that missing zero lol

jimthompson5910 (jim_thompson5910):

but I know what you mean

jimthompson5910 (jim_thompson5910):

so what's next?

OpenStudy (anonymous):

A= 9.04083087??

jimthompson5910 (jim_thompson5910):

not sure how you're getting that

OpenStudy (anonymous):

I did 3000^0.275

jimthompson5910 (jim_thompson5910):

ah i see now

jimthompson5910 (jim_thompson5910):

it should be 3000*e^0.275

jimthompson5910 (jim_thompson5910):

so you need to compute e^0.275 first

jimthompson5910 (jim_thompson5910):

the value of 'e' is roughly 2.71828, most likely there will be an 'e' button on your calculator so you won't have to worry about memorizing this

jimthompson5910 (jim_thompson5910):

anyways, e^0.275 = 2.71828^0.275 = 1.316530431 So 3000*e^0.275 turns into 3000*1.316530431

OpenStudy (anonymous):

Oh! I see =) So A= 3950

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

YES! =)

jimthompson5910 (jim_thompson5910):

you're on a roll lol